OpenStudy (anonymous):

A scale on a map shows that 2 inches = 25 miles How many inches on the map represent 60 miles?

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

is there a map?

OpenStudy (misty1212):

you want to solve \[\frac{2}{25}=\frac{x}{60}\] which you do by multiplying \[x=\frac{2\times 60}{25}\]

OpenStudy (anonymous):

A) 2 inches=25 miles x inches = 60 miles B) 2 inches=25 miles 1/4 inches= x miles

OpenStudy (anonymous):

so divide 25 by 2